How do i work out 5(x+4)=3x+28?
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

How do i work out 5(x+4)=3x+28?
--------------
Depends on what you mean by "work out."
If you want to solve for x (you should say that):
5(x+4)=3x+28?
5x+20 = 3x+28
2x = 8
x = 4

Answer by stanbon(75887)

5(x+4)=3x+28
-----------
5x + 20 = 3x + 28
-----
2x = 8
x = 4
--------------
Cheers,
Stan H.
==============